In triagle ABC, the measure of angle B is 41 degrees more than three times the measure of angle A. The measure of angle C is 64 degrees more than the measure of angle A. Find the measure of each angle.
Answer by jorel1380(3719)

a+b+c=180
b=3a+41
c=a+64
a+3a+41+a+64=180
5a=75
a=15
b=86
c=79..